OpenStudy (anonymous):

There are a total of 72 items. A. found 8 more than B. C found 2 times as many as B. What are the total items for each A, B and C found?

OpenStudy (anonymous):

Let A = number of items found by A, let B = number of items found by B, let C = number of items found by C, then A + B + C = 72 also A = B + 8 and C = 2B I suggest substituting, B + 8 for A in the first equation and C = 2B, also in the first equation.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."